OpenStudy (anonymous):

The rate law for the reaction 3A yields C is Rate=4.36E-2 L mol-1 hr-1 What is the half life for the reaction if the initial concentration of A is 0.250M?

OpenStudy (aaronq):

the rate law indicates that it's a zeroth order reaction, do you know which equations to use?

OpenStudy (anonymous):

I should be using the zeroth half-life order right? t1/2=0.5[A]0(k)

OpenStudy (aaronq):

yup it's divided by k though, not multiplied.

OpenStudy (anonymous):

t1/2 = 0.5(0.250M) / 4.36E-2 L x mol x hr?

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

My answer was 2.87

OpenStudy (aaronq):

cool

OpenStudy (anonymous):

But it doesn't match any of my answer that is giving though

OpenStudy (aaronq):

we'll that should be the correct answer. are the answers given in order of hours?

OpenStudy (anonymous):

Yeah the answers are given in hrs

OpenStudy (aaronq):

weird, because that IS the answer.

OpenStudy (anonymous):

Ok. Thanks

OpenStudy (aaronq):

no problem